OpenStudy (anonymous):

On the graph of the equation 3x + 2y = 18, what is the value of the y-intercept? –9 –6 6 9

OpenStudy (anonymous):

-9?

OpenStudy (anonymous):

it is 9

OpenStudy (anonymous):

how so?

OpenStudy (anonymous):

put x = 0 to find the Y intercept

OpenStudy (anonymous):

-9 works as well, i got x=12 which when plugged back into the formula, 18 equals 18

OpenStudy (anonymous):

3x + 2y = 18, put x = 0 0 + 2y = 18 y = 18/2 = 9

OpenStudy (anonymous):

oh but on a graph its positive not negative..okay okay thanks!

OpenStudy (anonymous):

 this is the graph